30 people from 10 families of 3 attend a meeting. If 5 people are asked to sit on 5 chairs on the stage and no two people come form the same family, how many arrangements are there?
Answer by Edwin McCravy(20054)

Choose a family for each chair in 10P5 = 10*9*8*7*6 ways.
Choose a family member to sit in each chairs in 3*3*3*3*3 = 35 ways.

Answer 10*9*8*7*6*35 = 7,348,320 ways.

Edwin